You have $410,000 invested in a well-diversified portfolio. You inherit a house that is presently worth $160,000. Consider the summary measures in the following table: Investment Expected Return Standard Deviation Old portfolio House 6% 13% 11% 19% The correlation coefficient between your portfolio and the house is o.5 a. What is the expected return and the standard deviation for your portfolio comprising your old portfolio and round intermediate calculations. Round your final answers to 2 decimal places.) the house? (Do not Expected return Standard deviation b. Suppose you decide to sell the house and use the proceeds of $160,000 to buy risk-free T-bills that promise a 8% rate of return. Calculate the expected return and the standard deviation for the resulting portfolio. [Hint. Note that the correlation coefficient between any asset and the risk-free T-bills is zero] (Do not round intermediate calculations. Round your final answers to 2 decimal places.) Expected return Standard deviation
Answer #1

here let x1 and x2 are return from old portfolio and House

a) therefore expected return =w1*E(x1)+w2*E(x2)=(41/57)*6+(16/57)*13 =7.96 %

standard deviation =sqrt((41*11/57)2+(16*19/57)2+2*0.5*(41/57)*(16/57)*11*19)=11.54%

b)

expected return =(41/57)*6+(16/57)*8 =6.56 %

standard deviation =sqrt((41*11/57)2+(16*0/57)2+2*0.0*(41/57)*(16/57)*11*0)=7.91%